# A comparison of pre-existing $\Lambda$CDM predictions with the abundance of JWST galaxies at high redshift

### v1

@@ Abstract
Observations with the James Webb Space Telescope have revealed a high abundance of bright galaxies at redshift, $z\gtrsim 12$, which has been widely interpreted as conflicting with the $\Lambda$CDM model. In Cowley et al. (2018) predictions were made – prior to the JWST observations – for the expected abundance of these galaxies using the Durham semi-analytic galaxy formation model, galform, which is known to produce a realistic population of galaxies at lower redshifts including the present day. Key to this model is the assumption of a “top-heavy” initial mass function of stars formed in bursts (required to explain the number counts and redshift distribution of sub-millimetre galaxies). Here, we compare the rest-frame ultraviolet luminosity functions derived from JWST observations with those predicted by the Cowley et al. model up to $z=14$ and make further predictions for $z=16$. We find that below $z\sim 10$, the Cowley et al. predictions agree very well with observations, while agreement at $z\gtrsim 12$ requires extending the model to take into account the timescale for the growth of obscuring dust grains and its dependence on gas metallicity. We trace the evolution of these galaxies from $z=14$ to $z=0$ and find that their descendants typically reside in halos with a median mass of $10^{13.6}\,h^{-1}\,\mathrm{M_{\odot}}$. The stellar masses of the descendants range from $10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$ to $10^{11.5}\,h^{-1}\,\mathrm{M_{\odot}}$. Although these galaxies were all central galaxies at $z=14$, nearly half of their descendants end up as satellites in massive halos.
@@ 1 Introduction
In the $\Lambda$CDM model, galaxies form within dark matter halos that evolve due to gravitational instability in the hierarchical way dictated by the amplitude and shape of the initial power spectrum of density perturbations (Frenk et al., 1985). Thus, the first galaxies form at high redshift in halos of relatively small mass which are, nevertheless, large enough for atomic hydrogen to cool within them (White & Rees, 1978; Blumenthal et al., 1984; White & Frenk, 1991; Benitez-Llambay & Frenk, 2020). The James Webb Space Telescope (JWST) was designed to search for these early galaxies. Detailed theoretical predictions for what the JWST was expected to find in deep galaxy surveys according to the $\Lambda$CDM model were published by Cowley et al. (2018) several years before the telescope was launched. That study presented the expected ultraviolet luminosity functions (UVLFs) in the JWST bands at high redshift, from $z=7$ to $z=16$, as calculated using the galform semi-analytic model of galaxy formation (Cole et al., 2000; Lacey et al., 2016). Other theoretical predictions followed, either using semi-analytic models (e.g. Yung et al., 2019) or cosmological hydrodynamical simulations (e.g. Ma et al., 2019; Vogelsberger et al., 2020).
@@
Very soon after the launch of JWST, a spate of papers was published claiming the detection of galaxies at very early times based on photometric redshifts. For example, Donnan et al. (2023) claimed to have discovered a galaxy at a record redshift of $z=16.4$ (CEERS-93316), only for the subsequent measurement of a spectrum to show that this galaxy is actually at $z=4.91$ (see also Zavala et al., 2023; Harikane et al., 2024). Similarly, Harikane et al. (2023) claimed to have found a galaxy at $z=16.41$ (S5-z16-1) which, however, as the authors themselves pointed out, has a spectral line detected by ALMA that would place it at $z=4.61$ (Fujimoto et al., 2023). Nevertheless, a number of the high redshift candidates have been spectroscopically confirmed to have $z>11$ (e.g. Arrabal Haro et al. 2023; Castellano et al. 2024; Hainline et al. 2024; Harikane et al. 2024).
@@
Several of these early observational studies published UVLFs at high redshift, based mostly on photometric redshifts (e.g. Bouwens et al., 2023; Donnan et al., 2023; Finkelstein et al., 2023; Harikane et al., 2023; Morishita & Stiavelli, 2023; Pérez-González et al., 2023; Robertson et al., 2023; Willott et al., 2023). Most of these papers claimed that the inferred UVLFs were inconsistent with the $\Lambda$CDM  model. For example, Finkelstein et al. (2023) say: “With this sample we confirm at higher confidence early JWST conclusions that bright galaxies in this epoch are more abundant than predicted by most theoretical models.” Similarly, Labbé et al. (2023) state that “… these stellar mass densities are difficult to realize in a standard $\Lambda$CDM cosmology …”. Unfortunately, all these papers overlooked the pre-existing $\Lambda$CDM predictions.
@@
Boylan-Kolchin (2023) published a simple calculation based on the mass function of dark matter halos which he compared to the stellar mass function inferred by Labbé et al. (2023). Boylan-Kolchin found that the baryons in halos turn into stars with some efficiency which has, implausibly, to be close to 1 (compared to $\sim 5\%-10\%$ in the Milky Way, e.g. Eke et al. 2005) in order to match the data, leading him to conclude that there is a tension with the $\Lambda$CDM model. This comparison is not trivial because it is the broadband spectral energy distribution (SED), not the stellar mass, that is observed, and inferring the stellar mass from the SED requires making assumptions that are highly uncertain (e.g. regarding the initial mass function and dust obscuration). A similar conclusion was reached from a similar calculation by Lovell et al. (2023) who stressed the limitations due to the uncertainties in the stellar masses and redshifts of the JWST galaxies. Indeed, Steinhardt et al. (2023) showed that using a new set of templates (including a varying IMF) to fit the SED, the inferred stellar masses of Labbé et al. (2023) are reduced by a factor of 10-50.
@@
Two studies have attempted to compare the JWST high redshift data with physically-based models in the $\Lambda$CDM cosmology. Kannan et al. (2023) used the $L=740~{}{\rm Mpc}$ MillenniumTNG hydrodynamic simulation (scaled to match galaxy properties from smaller-volume, higher-resolution simulations) to study the properties of galaxies at $z>10$, including UVLFs. They found broad agreement with the data up to $z=10$ but for $z=10-12$, they require the model galaxies to be dust-free and, beyond $z=12$, the model, even without dust, underpredicts the observed abundance by about one order of magnitude. They suggested that processes not included in the simulations, such as Population III stars, or a variable stellar initial mass function (IMF), may be required. Another physically-based model is that of Yung et al. (2024) based on the Santa Cruz semi-analytic model of galaxy formation (Somerville & Primack, 1999). They also predicted galaxy number densities at $z>11$ that are at least one order of magnitude below the observational data, even when they neglect dust obscuration. They also speculated that a top-heavy IMF might eliminate the discrepancy.
@@
Here, we compare the predictions of the Cowley et al. (2018) model, which implements galform in the large P-Millennium N-body simulation (Baugh et al., 2019), with the observed UVLFs from $z=7$ to $z=16$. We derive model scaling relations between fundamental properties such as star-formation rate or stellar mass as a function of UV luminosity and we calculate the fate and properties of the present-day descendants of these early JWST galaxies.
@@ 2 Method
2.1 The galform semi-analytic model and the P–Millennium N-body simulation
@@
Here, we follow the same procedure as in Cowley et al. (2018) and apply galform to an updated version of the Millennium simulation (Springel et al., 2005), which follows dark matter in a cube of sidelength $800\,\mathrm{Mpc}$, assuming cosmological parameter values derived from Planck (Planck Collaboration et al., 2016). We refer to this simulation as the P–Millennium; it was introduced by Baugh et al. (2019). The P–Millennium has a dark matter particle mass of $1.06\times 10^{8}\,h^{-1}\,\mathrm{M_{\odot}}$, with halos defined initially with a friends-of-friends algorithm (Davis et al., 1985), in which self-bound structures are identified with the subfind algorithm (Springel et al., 2001). Merger trees are constructed from these subhalo catalogues using the ‘DHalos’ algorithm (Jiang et al., 2014). The large volume and high mass resolution of P-Millennium allow us not only to resolve halos down to $\sim 2\times 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$ (20 dark matter particles), but also to study the rare massive halos that may be the hosts of bright galaxies at high redshift, $z\sim 14-16$. In Fig. 1, we show the halo mass functions in P–Millennium from $z=7$ to $z=16$. We observe that even at $z=16$, halos as massive as $10^{10.3}\,h^{-1}\,\mathrm{M_{\odot}}$ have a number density as high as $\sim 10^{-7}\,h^{3}\,\mathrm{Mpc^{-3}}$, which corresponds to 16 halos/dex in the simulation box.
Throughout this paper, we assume a flat universe and, to be consistent with the P–Millennium, we adopt cosmological parameters from Planck Collaboration et al. (2016, i.e. $\Omega_{\rm m}=0.307$, $\Omega_{\Lambda}=0.693$, $\Omega_{\rm b}=0.0483$, $\sigma_{8}=0.8288$, and $h=0.6777$).
Figure 1: Halo mass functions measured in the P–Millennium N-body simulation between $z=7$ and $z=16$. We use the galform semi-analytic model to predict galaxy populations at each of these redshifts.
@@ 2.1.1 Supernova feedback model
where $\tau_{\text{dyn,halo }}$ is the halo dynamical time and $\alpha_{\text{ret }}$ is a constant, set here to $\alpha_{\text{ret }}=1.0$. We refer to this as the standard or STD feedback model.
A deficiency of the STD model is that it does not produce enough ionizing photons at early times to reionize the Universe at the high redshift implied by the Planck Collaboration et al. (2016) data. This could be remedied by reducing the strength of supernova feedback at high redshift, leading to enhanced star formation at early times. One consequence of this is that, with the original parametrization of feedback, this would result in the overproduction of dwarf galaxies today as well as in an incorrect stellar mass-metallicity relation for these galaxies. This led Hou et al. (2016) to propose a new SN feedback model for galform in which the feedback strength changes with cosmic time. In this updated model, which we refer to as EvolFB, the simple power-law “mass loading” factor, $\beta(V_{\rm c})$ (Eq. missing 1) is replaced by a broken power law as a function of $V_{\rm c}$:
@@ 2.1.2 Stellar initial mass function
where $N$ is the number of stars as a function of their mass at birth, $m$. In galform we distinguish between two modes of star formation (SF), each with a different IMF: (i) a quiescent (or disc) SF with a Kennicutt (1983) IMF ($x=0.4$ for $m<1\,\mathrm{M_{\odot}}$ and $x=1.5$ for $m>1\,\mathrm{M_{\odot}}$), and (ii) a starburst mode with a top-heavy IMF ($x=1$ for the entire mass range, $0.1\,\mathrm{M_{\odot}}$ to $100\,\mathrm{M_{\odot}}$). As we mentioned in the Introduction (Section 1), Baugh et al. (2005) found that a top-heavy IMF in starbursts is necessary to account for the number counts and redshift distribution of sub-millimetre galaxies (SMGs), while remaining consistent with observational constraints at $z=0$ (see also Lacey et al. 2016).
Since the frequency of starbursts is higher at higher redshift, the dual IMF in galform is, in practice, equivalent to a redshift-dependent IMF which, on average, becomes increasingly top-heavy at higher redshift. We note here that although a redshift dependence of the IMF remains an open question, more and more evidence has accumulated suggesting the presence of a top-heavy IMF in the early Universe. For example, Sneppen et al. (2022) fitted a temperature-dependent IMF to galaxies in the COSMOS2015 catalogue (Laigle et al., 2016) and found that the IMF becomes progressively top-heavy with increasing redshift. Cameron et al. (2023) also reported two Lyman-$\alpha$ emitting galaxies at redshifts 5.9 and 7.9 that show evidence for exceptionally top-heavy IMFs.

### v2

@@ Abstract
Observations with the James Webb Space Telescope have revealed a high abundance of bright galaxies at redshift, $z\gtrsim 12$, which has been widely interpreted as conflicting with the $\Lambda$CDM model. In Cowley et al. (2018) predictions were made – prior to the JWST observations – for the expected abundance of these galaxies using the Durham semi-analytic galaxy formation model, galform, which is known to produce a realistic population of galaxies at lower redshifts including the present day. Key to this model is the assumption of a “top-heavy” initial mass function of stars formed in bursts (required to explain the number counts and redshift distribution of sub-millimetre galaxies). Here, we compare the rest-frame ultraviolet luminosity functions derived from JWST observations with those predicted by the Cowley et al. model up to $z=14$ and make further predictions for $z=16$. We find that below $z\sim 10$, the Cowley et al. predictions agree very well with observations, while agreement at $z\gtrsim 12$ requires extending the model to take into account the timescale for the growth of obscuring dust grains at these very early times and its dependence on gas metallicity. We trace the evolution of these galaxies from $z=14$ to $z=0$ and find that their descendants typically reside in halos with a median mass $2.5\times 10^{13}\,h^{-1}\,\mathrm{M_{\odot}}$. The stellar masses of the descendants range from $3.2\times 10^{6}\,h^{-1}\,\mathrm{M_{\odot}}$ to $3.2\times 10^{11}\,h^{-1}\,\mathrm{M_{\odot}}$. Although these galaxies were all central galaxies at $z=14$, over half of their descendants end up as satellites in massive halos.
@@ 1 Introduction
In the $\Lambda$CDM model, galaxies form within dark matter halos that evolve due to gravitational instability in the hierarchical way dictated by the initial power spectrum of density perturbations (Frenk et al., 1985). Thus, the first galaxies form at high redshift in halos of relatively small mass which are, nevertheless, large enough for atomic hydrogen to cool within them (White & Rees, 1978; Blumenthal et al., 1984; White & Frenk, 1991; Benitez-Llambay & Frenk, 2020). The James Webb Space Telescope (JWST) was designed to search for these early galaxies. Detailed theoretical predictions for what the JWST was expected to find in deep galaxy surveys according to the $\Lambda$CDM model were published by Cowley et al. (2018) several years before the telescope was launched. That study presented the expected ultraviolet luminosity functions (UVLFs) in the JWST bands at high redshift, from $z=7$ to $z=16$, as calculated using the galform semi-analytic model of galaxy formation (Cole et al., 2000; Lacey et al., 2016). Other predictions followed, using semi-analytic models (e.g. Yung et al., 2019) or cosmological hydrodynamical simulations (e.g. Ma et al., 2019; Vogelsberger et al., 2020).
@@
Soon after the launch of JWST, a spate of papers was published claiming the detection of galaxies at very early times based on photometric redshifts. For example, Donnan et al. (2023) claimed to have discovered a galaxy at a record redshift of $z=16.4$ (CEERS-93316), only for the subsequent measurement of a spectrum to show that this galaxy is actually at $z=4.91$ (see also Zavala et al., 2023; Harikane et al., 2024). Similarly, Harikane et al. (2023) claimed to have found a galaxy at $z=16.41$ (S5-z16-1) which, however, as the authors themselves pointed out, has a spectral line detected by ALMA, placing it at $z=4.61$ (Fujimoto et al., 2023). Nevertheless, a number of the high redshift candidates have been spectroscopically confirmed to have $z>11$ (e.g. Arrabal Haro et al. 2023; Castellano et al. 2024; Hainline et al. 2024; Harikane et al. 2024; Zavala et al. 2024).
@@
Several of these early studies published UVLFs at high redshift, based mostly on photometric redshifts (e.g. Bouwens et al., 2023; Donnan et al., 2023; Finkelstein et al., 2023; Harikane et al., 2023; Morishita & Stiavelli, 2023; Pérez-González et al., 2023; Robertson et al., 2023; Willott et al., 2023). Most of these papers claimed that the inferred UVLFs were inconsistent with $\Lambda$CDM. For example, Finkelstein et al. (2023) say: “With this sample we confirm at higher confidence early JWST conclusions that bright galaxies in this epoch are more abundant than predicted by most theoretical models.” Similarly, Labbé et al. (2023) state that “… these stellar mass densities are difficult to realize in a standard $\Lambda$CDM cosmology …”. Unfortunately, all these papers overlooked the pre-existing $\Lambda$CDM predictions.
@@
Boylan-Kolchin (2023) published a simple calculation based on the mass function of dark matter halos which he compared to the stellar mass function inferred by Labbé et al. (2023). Boylan-Kolchin found that the baryons in halos would have turn into stars with an efficiency which has, implausibly, to be close to 1 (compared to $\sim 5\%-10\%$ in the Milky Way, e.g. Eke et al. 2005) in order to match the data, leading him to conclude that there is a tension with the $\Lambda$CDM model. This comparison is not trivial because it is the broadband spectral energy distribution (SED), not the stellar mass, that is observed, and inferring the stellar mass from the SED requires making assumptions that are highly uncertain (e.g. regarding the initial mass function and dust obscuration). A similar conclusion was reached from a similar calculation by Lovell et al. (2023) who stressed the limitations due to the uncertainties in the stellar masses and redshifts of the JWST galaxies. Indeed, Steinhardt et al. (2023) showed that using a new set of templates (including a varying IMF) to fit the SED, the inferred stellar masses of Labbé et al. (2023) are reduced by a factor of 10-50. Chworowsky et al. (2023) analyzed JWST data for a sample of galaxies at $z=4-8$ and, using a similar argument to that of Boylan-Kolchin (2023), concluded that while the abundance of galaxies at $z\sim 5$ could be accounted for within $\Lambda$CDM assuming a conventional efficiency of baryon conversion into stars ($\sim 0.14$), the efficiency would again have to become implausibly large at higher $z$ unless their inferred stellar masses had been significantly overestimated.
@@
Two studies have attempted to compare the JWST high redshift data with physically-based models in the $\Lambda$CDM cosmology. Kannan et al. (2023) used the $L=740~{}{\rm Mpc}$ MillenniumTNG hydrodynamic simulation (scaled to match galaxy properties from smaller-volume, higher-resolution simulations) to study the properties of galaxies at $z>10$, including UVLFs. They found broad agreement with the data up to $z=10$ but for $z=10-12$, they require the model galaxies to be dust-free and, beyond $z=12$, the model, even without dust, underpredicts the observed abundance by about one order of magnitude. They suggested that processes not included in the simulation, such as Population III stars, or a variable stellar initial mass function (IMF), may be required. Another physically-based model is that of Yung et al. (2024) based on the Santa Cruz semi-analytic model (Somerville & Primack, 1999). They also predicted galaxy number densities at $z>11$ that are at least one order of magnitude below the observational data, even when they neglect dust obscuration. They also speculated that a top-heavy IMF might ease the discrepancy.
@@
Here, we compare the predictions of the Cowley et al. (2018) model with the observed UVLFs from $z=7$ to $z=16$. We derive model scaling relations between fundamental properties such as star formation rate, stellar mass, halo mass, and UV luminosity. We also calculate the fate and properties of the present-day descendants of these early JWST galaxies.
@@ 2 Method
2.1 The galform semi-analytic model
@@
2.1.1 The dark matter-only simulations
Cowley et al. (2018) applied galform to the P–Millennium simulation, a variant of the Millennium simulation (Springel et al., 2005), with cosmological parameter values derived from Planck data (Planck Collaboration et al., 2016), in a cube of sidelength $800\,\mathrm{Mpc}$. The P–Millennium simulation was introduced by Baugh et al. (2019). Halos were initially identifed with a friends-of-friends algorithm (Davis et al., 1985) and self-bound structures were then identified with the subfind algorithm (Springel et al., 2001). Merger trees were constructed from these subhalo catalogues using the ‘DHalos’ algorithm (Jiang et al., 2014).
The large volume of P–Millennium allows us to study the rare massive halos that could host bright galaxies at high redshift, $z\sim 14-16$, but the relatively large dark matter particle mass, $1.06\times 10^{8}\,h^{-1}\,\mathrm{M_{\odot}}$, limits the ability to follow the low-mass halos that could host very faint galaxies. To overcome this resolution limitation, we combine the P–Millennium with the EAGLE–DMO (dark matter-only) simulation (Guo et al., 2016). EAGLE–DMO assumed the same cosmological parameters as the P–Millennium, but it has much higher mass resolution, a particle mass of $1.15\times 10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$, 10 times smaller than P–Millennium. To combine the two simulations, galaxies with halo masses larger than $5\times 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$ (roughly corresponding to 50 DM particles in P–Millennium and 450 DM particles in EAGLE–DMO) are taken from P–Millennium, while those with lower halo masses are taken from the EAGLE–DMO simulation.
Since the EAGLE–DMO simulation has a cube sidelength of $100\,{\rm Mpc}$, $1/8$ that of P–Millennium, when calculating the luminosity functions we assign a weight $5128^{3}=$ 512 larger to the EAGLE–DMO galaxies than to the P–Millennium galaxies. We will return to this point in Section 3. With this combination, we are able not only to study high redshift rare massive objects, but also low-mass halos down to $\sim 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$ (about 100 DM particles in EAGLE–DMO).
For P–Millennium, there are 271 snapshots in total, 15 between $z=12$ and $z=16$; for EAGLE–DMO, there are 400 snapshots in total, 14 between $z=12$ and $z=16$. This allows us accurately to capture the merging history of halos and galaxies. In Fig. 1, we show the halo mass functions in both P–Millennium (solid) and EAGLE–DMO (dashed) from $z=7$ to $z=16$. We observe that even at $z=16$, halos as massive as $2\times 10^{10}\,h^{-1}\,\mathrm{M_{\odot}}$ have a number density as high as $\sim 10^{-7}\,h^{3}\,\mathrm{Mpc^{-3}}$, which corresponds to 16 halos/dex in the simulation cube.
Throughout we assume a flat universe and, to be consistent with the P–Millennium and EAGLE–DMO, we adopt cosmological parameters from Planck Collaboration et al. (2016, i.e. $\Omega_{\rm m}=0.307$, $\Omega_{\Lambda}=0.693$, $\Omega_{\rm b}=0.0483$, $\sigma_{8}=0.8288$, and $h=0.6777$).
Figure 1: Halo mass functions measured in the P–Millennium (solid) and EAGLE–DMO (dashed) N-body simulations between $z=7$ and $z=16$. The vertical dashed-dotted line indicates $M_{\rm halo}=5\times 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$, which corresponds to the mass of 50 particles in P–Millennium.
@@ 2.1.2 Supernova feedback model
where $\tau_{\text{dyn,halo }}$ is the halo dynamical time and $\alpha_{\text{ret }}$ is a constant, set to $\alpha_{\text{ret }}=1.0$. We refer to this as the standard (STD) feedback model.
A deficiency of the STD model is that it does not produce enough ionizing photons at early times to reionize the Universe at the high redshift implied by the Planck Collaboration et al. (2016) data. This could be remedied by reducing the strength of supernova feedback at high redshift, leading to enhanced star formation at early times. One consequence of this is that, with the original parametrization of feedback, this would result in the overproduction of dwarf galaxies today as well as in an incorrect stellar mass-metallicity relation for these galaxies. This led Hou et al. (2016) to propose a new SN feedback model for galform in which the feedback strength changes with cosmic time. In this model, referred to as EvolFB, the simple power-law “mass loading” factor, $\beta(V_{\rm c})$ (Eq. missing 1) is replaced by a broken power law as a function of $V_{\rm c}$:
@@ 2.1.3 Stellar initial mass function
where $N$ is the number of stars as a function of their mass at birth, $m$. In galform we distinguish between two modes of star formation (SF): (i) a quiescent (or disc) SF with a Kennicutt (1983) IMF ($x=0.4$ for $m<1\,\mathrm{M_{\odot}}$ and $x=1.5$ for $m>1\,\mathrm{M_{\odot}}$), and (ii) a starburst mode with a top-heavy IMF ($x=1$ for the entire mass range, $0.1\,\mathrm{M_{\odot}}$ to $100\,\mathrm{M_{\odot}}$). As we mentioned in the Introduction (Section 1), Baugh et al. (2005) found that a top-heavy IMF in starbursts is necessary to account for the number counts and redshift distribution of sub-millimetre galaxies (SMGs), while remaining consistent with observational constraints at $z=0$ (see also Lacey et al. 2016).
Since the frequency of starbursts is higher at higher redshift, the dual IMF in galform is, in practice, equivalent to a redshift-dependent IMF which, on average, becomes increasingly top-heavy at higher redshift. We note that although a redshift dependence of the IMF remains an open question, evidence has accumulated suggesting the presence of a top-heavy IMF in the early Universe. For example, Sneppen et al. (2022) fitted a temperature-dependent IMF to galaxies in the COSMOS2015 catalogue (Laigle et al., 2016) and found that the IMF becomes progressively top-heavy with increasing redshift. Cameron et al. (2023) also reported two Lyman-$\alpha$ emitting galaxies at redshifts 5.9 and 7.9 that show evidence for exceptionally top-heavy IMFs. Liang et al. (2021) found that the IMF tends to be more top-heavy at lower metallicities, supporting the possibility of a top-heavy IMF at high redshift.
2.1.4 Stellar population and luminosity
The stellar population synthesis model in galform is explained in Lacey et al. (2016). Here, we provide a brief overview of the procedure for deriving the properties of the stellar populations and the luminosities in different bands for galaxies in galform.
To obtain galaxy luminosities, galform calculates spectral energy distributions (SEDs) from a set of single stellar population (SSP) templates. The SED of a stellar population at a given time, $t$, can be written as:
[…]
where $\Psi(t^{\prime},Z^{\prime})\,dt^{\prime}\,dZ^{\prime}$ represents the mass of stars at birth which formed in the time interval $(t^{\prime},t^{\prime}+dt^{\prime})$ and metallicity range $(Z^{\prime},Z^{\prime}+dZ^{\prime})$. $L^{\rm(SSP)}_{\lambda}(t,Z;\Phi)$ is the SED of an SSP of one solar mass with age $t$ and metallicity $Z$, given a specific IMF, $\Phi(m)$. $\Psi(t,Z)$ is derived by summing over the star formation histories of all the progenitor galaxies which merged into the final galaxy. The SSP luminosity is determined by integrating the luminosity, $L^{\rm(star)}(t,Z,m)$, of a star of initial mass $m$, metallicity $Z$, and age $t$ over the IMF:
$L^{\rm(SSP)}_{\lambda}(t,Z;\Phi)=\int_{m_{L}}^{m_{U}}\,L^{\rm(star)}_{\lambda}(t,Z,m)\,\Phi(m)\,d\ln m,$ (7)
[…]